1.Write as decimals, correct to two decimal places
91/16
2.find 3% of 200 pesos
3.a municipality levies a 12% sales tax. Find the tax on a car listed 350,000 pesos

Answer by Boreal(15235)

91/16=5 and 11/16
11/16=0.6875
5.69
200*0.03=6 3% of 100 is 3, and 3% of 200 is 6.
350,000*0.12=42,000
Ten per cent would be 35,000
one per cent is 3500
two per cent is 7000
Twelve per cent is 42000